OpenStudy (anonymous):

7 square root of 3 end root minus 4 square root of 6 end root plus square root of 48 end root minus square root of 54

OpenStudy (anonymous):

would it be: 11 square root of 6 end root minus 7 square root of 12 11 square root of 3 end root minus 7 square root of 6 negative 3 square root of 9 4 square root of 9

OpenStudy (anonymous):

\[7\sqrt{3}-4\sqrt{6}+\sqrt{48}-\sqrt{54}\] Is the first part right?

OpenStudy (imstuck):

The goal is to get all the radicands to be either a sqrt 3 or a sqrt 6 so you can combine the like radicands. simplify the sqt 48 and the sqrt 54.

OpenStudy (anonymous):

@Cosmichaotic yes that is right

OpenStudy (anonymous):

@IMStuck ok

OpenStudy (anonymous):

@IMStuck i would use 6x8 and 6x9